OpenStudy (anonymous):

Use MI to prove that \(n^2-1\) is divisible by 8 for any positive odd integer n.

OpenStudy (anonymous):

Attempt: Let p(n) be the statement \(n^2−1\) is divisible by 8 When n = 1 p(1) = 1^2 - 1 = 0, which is divisible by 8. So, p(1) is true. Suppose p(k) is true for some positive odd integers k, i.e. k^2-1 = 8m, where \(m\in\mathbb{Z^+}\) Consider p(k+2) \[(k+2)^2 - 1\]\[=k^2+4k+4-1\]\[=8m+1+4k+4-1\]\[=8m+4(k+1)\] Then, I don't know how I should proceed from here.

OpenStudy (anonymous):

P(n)=n^2-1,where n is an odd integer. (1) Let n=2k+1,where k is a positive integer. P(n)=〖(2k+1)〗^2-1 ….(2) P(1)=3^2-1=8,which is divisible by 8 Assume that p(m) is true. P(m)=〖(2m+1)〗^2-1 is divisible by 8 Therefore let (2m+1)^2-1=8l (2m+1)^2=8l+1 P(m+1)={2(m+1)+1}^2-1={(2m+1)+2}^2-1 =(2m+1)^2+4+2*2(2m+1)-1 =(2m+1)^2+4+8m+4-1 =8l+1+8+8m-1 =8 (l+m+1) =8*an integer Hence (2) is true for all k Or (1) is true for all positive odd integers.

OpenStudy (anonymous):

"Let n=2k+1,where k is a positive integer" What about the case when n=1?

OpenStudy (anonymous):

let m^2-1=8l ,where l is an odd integer. m^2=8l+1 Next odd integer=m+2 P(m+2)=(m+2)^2-1=m^2+4m+4-1 =m^2 +4(m+1)-1 =8l+1+4(m+1)-1 =8l+4(m+1) =8l+8t, m is odd so m+1 is even ,or m+1=2t (say) =8(l+t) =8 * an integer Hence p(m+2) is true. Hence by induction p(n) is true for all positive integers

OpenStudy (anonymous):

correction i should have written k=0,1,2,3,............ and in the beginning p(0)

OpenStudy (anonymous):

Ah! I am stupid :\ 4(m+1) = 8t is the key!

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

yw